in a heating experiment is was noted that temp of liquid in beaker rose at 4k per min just before it began to boil and that 40 min later all liquid had boiled away . for this liquid what is numerical ratio( specific heat capacity /specific latent heat of vaporization? 
ans is 1/160

Grade:12th pass

1 Answers

Rituraj Tiwari
3 years ago
Temperature of the liquid after 40 minutes = 4 * 40 = 160K

Specific latent heat of vaporization = 160

Specific heat capacity of the liquid = 1

Therefore,numerical ratio = specific heat capacity /specific latent heat of vaporization = 1 / 160